# A Comprehensive Survey of Hallucination Mitigation Techniques in Large Language Models

### 02 Jan 2024 v1

@@ A Comprehensive Survey of Hallucination Mitigation Techniques in Large Language Models
S.M Towhidul Islam Tonmoy1, S M Mehedi Zaman1, Vinija Jain3,4*, Anku Rani2, Vipula Rawte2, Aman Chadha3,4thanks:    Work does not relate to position at Amazon.  , Amitava Das2 1Islamic University of Technology, Bangladesh 2AI Institute, University of South Carolina, USA 3Stanford University, USA, 4Amazon AI, USA towhidulislam@iut-dhaka.edu
@@ 2.1.2 Self-refinement through feedback and reasoning
Mind’s Mirror: While chain-of-thought (CoT) distillation methods like CoT distillation show promise in downsizing LLMs to small language models (SLMs), there is a risk of carrying over flawed reasoning and hallucinations. To address this, Liu et al. (2023) proposed methodology has two key components: firstly, a novel approach introduces the distillation of the self-evaluation capability inherent in LLMs into SLMs, aiming to mitigate adverse effects and reduce hallucinations. Secondly, a comprehensive distillation process incorporates multiple distinct CoT and self-evaluation paradigms for a holistic knowledge transfer into SLMs. The innovative methodology trains SLMs to possess self-evaluation capabilities, recognizing and correcting hallucinations and unreliable reasoning, enhancing predictive accuracy and reliability in various NLP tasks. To ensure comprehensive thinking, the methodology involves distilling SLMs from diverse CoTs and self-evaluation outputs generated by LLMs, providing a broader range of capabilities. Comprehensive experiments demonstrate the superiority of this method across varied reasoning tasks, offering a promising approach to downsizing LLMs for practical deployment.
@@
DRESS: Chen et al. (2023) propose the distinct use of natural language feedback (NLF), specifically categorized into critique and refinement NLF, to improve the alignment with human preferences and interaction capabilities of large vision language model (LVLM) Gan et al. (2020). They generalize the conditional reinforcement learning algorithm to effectively incorporate the NLF, which is non-differentiable, by training the model to generate corresponding responses conditioned on the NLF. They systematically evaluate the proposed LVLM, DRESS, on helpfulness, honesty, and harmlessness alignment and show relative improvements of 9.76%, 11.52%, and 21.03% compared to the prior state-of-the-art.
@@
MixAlign: Despite having accurate reference points, LLMs may choose to disregard them and instead rely on incorrect references or their own biases. This tendency to hallucinate arises particularly when users, lacking detailed knowledge of stored information, ask questions that may not align directly with the retrieved reference points. Zhang et al. (2023b) focuses on this knowledge alignment problem and introduces MixAlign, a framework that interacts with both the human user and the knowledge base to obtain and integrate clarifications on how the user question relates to the stored information. MixAlign employs a language model to achieve automatic knowledge alignment and, if necessary, further enhances this alignment through human user clarifications. MixAlign focuses on enhancing its ability to utilize grounding knowledge for faithful decision-making and issue resolution. In cases where the mapping process yields uncertainties or the evidence remains unclear, MixAlign generates a question seeking further clarification from the user, a step we refer to as human-assisted knowledge alignment.
Chain-of-Verification (CoVe): Dhuliawala et al. (2023) develop the CoVe method whereby the model first (i) drafts an initial response; then (ii) plans verification questions to fact-check its draft; (iii) answers those questions independently so the answers are not biased by other responses; and (iv) generates its final verified response. In experiments, they show that CoVe decreases hallucinations across a variety of tasks, from list-based questions from Wikidata closed book MultiSpanQA and long-form text generation. Given a user query, an LLM generates a baseline response that may contain inaccuracies, e.g., factual hallucinations. They show a query that failed for GPT-3.5. To improve this, CoVe first generates a plan of a set of verification questions to ask and then executes that plan by answering them and hence checking for agreement. They find that individual verification questions are typically answered with higher accuracy than the original accuracy of the facts in the original long-form generation. Finally, the revised response takes into account the verifications. The factored version of CoVe answers verification questions such that they cannot condition on the original response, avoiding repetition and improving performance. CoVe also uses a related self-consistency approach, but without the multi-agent (multi-LLM) debate concept. Their overall process, which they call CoVe, thus performs four core steps: 1. Generate Baseline Response: Given a query, generate the response using the LLM. 2. Plan Verifications: Given both query and baseline response, generate a list of verification questions that could help to self-analyze if there are any mistakes in the original response. 3. Execute Verifications: Answer each verification question in turn, and hence check the answer against the original response to check for inconsistencies or mistakes. 4. Generate Final Verified Response: Given the discovered inconsistencies (if any), generate a revised response incorporating the verification results. They note that verification questions are not templated and the LLM is free to phrase these in any form they want, and they also do not have to closely match the phrasing of the original text.
In their experiments, they perform such verification planning by providing a few-shot prompt of (response, verification) demonstrations to their LLM. While techniques such as retrieval-augmentation could be used in this process, such as verification via a search engine, this work does not explore tool use.
They consider only using the LLM itself in all steps of CoVe, hence the model is used to check its work. They investigate several variants of verification execution, called joint, 2-step, factored, and factor+revise.
@@
Chain of Natural Language Inference (CoNLI): Lei et al. (2023) address the challenge of hallucinations generated by LLMs when provided with background context. Despite their fluency in generating natural language, LLMs often produce ungrounded hallucinations not supported by the given sources. The proposed hierarchical framework focuses on detecting and mitigating such hallucinations without requiring fine-tuning or domain-specific prompts. The framework utilizes CoNLI for hallucination detection and achieves state-of-the-art performance in identifying ungrounded content. The approach involved post-editing to reduce hallucinations and enhance text quality without additional model adjustments. Through extensive experiments on various text-to-text datasets, the framework proves effective in both hallucination detection and reduction. By formulating hallucination detection as a chain of Natural Language Inference tasks, the framework incorporates both sentence-level and entity-level judgments, demonstrating interpretability in its output. The generalizability of this plug-and-play framework allows for seamless deployment across different contexts, showcasing competitive performance in hallucination detection and reduction while preserving text quality.
@@ 2.1.3 Prompt Tuning
SynTra: LLMs often exhibit hallucination in abstractive summarization tasks, even when all necessary information is present in the context. Addressing this challenge is difficult due to the intricate evaluation of hallucination at each optimization step. Jones et al. (2023) introduce SynTra, a method that leverages a synthetic task to efficiently reduce hallucination on real-world downstream tasks. SynTra optimizes the LLM’s system message through prefix-tuning on the synthetic task and transfers this enhanced message to more challenging, realistic tasks. The study demonstrates a reduction in hallucination for two 13B-parameter LLMs across three abstractive summarization tasks, highlighting the effectiveness of working with synthetic data to mitigate undesired behaviors. While SynTra proves successful, there are potential improvements to explore. Refinements in the optimization method, such as searching for better hyperparameters and exploring alternative fine-tuning methods, could enhance performance. Adapting SynTra for models accessible only via APIs and scaling it to larger open-source models are also avenues for further exploration. The study acknowledges the need for a more systematic approach in defining synthetic tasks and understanding the causal properties that reduce hallucination. The work raises important questions about the effectiveness of synthetic tasks compared to collecting human feedback directly on realistic tasks. Synthetic tasks offer scalability and clarity in biases but require a loss-inducing transfer step. Understanding when synthetic data is more effective than demonstrations could contribute to better resource allocation in ensuring the reliability, helpfulness, and safety of LLMs.
@@ 3.3 Introducing faithfulness based loss function
Creating a metric to gauge how closely a model’s outputs match input data or ground truth is the task of this section. In this sense, faithfulness describes the model’s capacity to faithfully and properly reflect data from the input without adding errors, omissions, or distortions Chrysostomou and Aletras (2021). The following methods portray the use of technique: Text Hallucination Mitigating (THAM) Framework: Yoon et al. (2022) introduce the THAM framework for Video-grounded Dialogue. THAM considers the text hallucination problem, which copies input texts for answer generation without the understanding of the question. It mitigates feature-level hallucination effects by introducing information-theoretic regularization. THAM framework incorporates Text Hallucination Regularization (THR) loss derived from the mutual information between the response language model and the proposed hallucination language model. Minimizing THR loss contributes to reducing indiscriminate text copying and boosting dialogue performances. THAM framework incorporates Text Hallucination Regularization loss derived from the proposed information-theoretic text hallucination measurement approach.
@@ 3.4 Supervised fine-tuning (SFT)
Knowledge Injection and Teacher-Student Approaches: Elaraby et al. (2023) focuses on measuring and reducing hallucinations in BLOOM 7B Workshop et al. (2022), a representative of such weaker open-source LLMs that are publicly available for research and commercial applications. They introduce HALOCHECK, a lightweight BlackBox knowledge-free frame- work designed to quantify the severity of hallucinations in LLMs. They explore techniques like knowledge injection and teacher-student approaches to alleviate hallucinations in low-parameter LLMs. The framework employs sentence-level entailment techniques to quantitatively assess the extent of hallucination within the generated output. The work aims to enhance the knowledge of smaller LLMs through Knowledge Injection (KI). This method involves fine-tuning the small LLM with domain-specific knowledge, without relying on manual instructions or instructions obtained from stronger LLMs, which can be resource-intensive to generate at scale. The HALO framework is illustrated in three parts. (a) HALOCHECK framework components for hallucinations estimation. (b) Knowledge injection training aims to reduce hallucinations by injecting abstract knowledge representations. (c) Improving student responses by stronger teacher responses, wherein teacher involvement is selectively triggered to manage the frequency of teacher involvement. They investigate leveraging a more powerful LLM, specifically focusing on GPT-4, to guide weaker LLMs by generating detailed answers to questions. Through assessing the severity of hallucination, we aim to optimize the engagement of the teacher LLM, thereby reducing the need for frequent queries to the teacher model. This strategy offers the advantage of alleviating the computational load and costs associated with heavy reliance on extensive-parameter LLMs.
@@
Refusal-Aware Instruction Tuning (R-Tuning): Zhang et al. (2023a) present a new approach called R-Tuning. This approach is formalized by first identifying the knowledge gap between parametric knowledge and the instruction tuning data. Then, we construct the refusal-aware data based on the knowledge intersection, to tune LLMs to refrain from responding to questions beyond its parametric knowledge. R-Tuning aims to endow the model with refusal-aware answering ability by recognizing when they should and shouldn’t claim knowledge. R-Tuning introduces two steps:
@@
    Measure the knowledge gap between parametric knowledge and the instruction tuning data and identify uncertain questions. By inferring the model on the training data once and comparing the prediction and label, the instruction tuning data is split into uncertain data D0 and certain data D1.
@@
    Construct the refusal-aware data by padding the uncertainty expression after the label words and then finetuning the model on the refusal-aware data.
Think While Effectively Articulating Knowledge (TWEAK): To reduce hallucinations, Qiu et al. (2023a) propose a novel decoding method, TWEAK. TWEAK treats the generated sequences at each decoding step and its future sequences as hypotheses, and ranks each generation candidate based on how well their corresponding hypotheses support the input facts using a Hypothesis Verification Model (HVM). The authors tweak only the decoding process without requiring re-training of the generative models, thus making our approach easily integrated with any Knowledge-to-text generation (K2T) generator. The existing decoding methods of a generative model, such as beam search, sample candidates only from the predicted likelihood, without any consideration of the faithfulness implication of these candidates. The authors propose a new dataset, FATE, which pairs and aligns input facts with their original and counterfactual descriptions at the word level.

### 03 Jan 2024 v2

@@ A Comprehensive Survey of Hallucination Mitigation Techniques in Large Language Models
S.M Towhidul Islam Tonmoy1, S M Mehedi Zaman1, Vinija Jain3,411footnotemark: 1, Anku Rani2, Vipula Rawte2, Aman Chadha3,4thanks:    Work does not relate to position at Amazon.  , Amitava Das2 1Islamic University of Technology, Bangladesh 2AI Institute, University of South Carolina, USA 3Stanford University, USA, 4Amazon AI, USA towhidulislam@iut-dhaka.edu
@@ 2.1.2 Self-refinement through feedback and reasoning
Mind’s Mirror: While chain-of-thought (CoT) distillation methods show promise for downsizing LLMs to small language models (SLMs), there is a risk of carrying over flawed reasoning and hallucinations. To address this, Liu et al. (2023) proposed a methodology with two key components: First, a novel approach introduces distilling the self-evaluation capability inherent in LLMs into SLMs, aiming to mitigate adverse effects and reduce hallucinations. Second, a comprehensive distillation process incorporates multiple distinct CoT and self-evaluation paradigms for holistic knowledge transfer into SLMs.
The methodology trains SLMs to possess self-evaluation capabilities, recognizing and correcting hallucinations and unreliable reasoning, enhancing predictive accuracy and reliability on various NLP tasks. Comprehensive experiments demonstrate the superiority of this method across reasoning tasks, offering a promising approach to responsibly downsize LLMs.
@@
DRESS: Chen et al. (2023) propose using natural language feedback (NLF), specifically critique and refinement NLF, to improve alignment with human preferences and interaction capabilities of large vision language models (LVLMs). They generalize conditional reinforcement learning to effectively incorporate non-differentiable NLF by training the model to generate corresponding responses conditioned on the NLF. Experiments show relative improvements in DRESS over prior state-of-the-art LVLMs in metrics of helpfulness, honesty, and harmlessness alignment.
@@
MixAlign: Despite having accurate reference points, LLMs may disregard them and rely on incorrect references or biases instead. This tendency to hallucinate arises when users ask questions that do not directly align with the retrieved references, lacking detailed knowledge of the stored information. Zhang et al. (2023b) focus on this knowledge alignment problem and introduce MixAlign, a framework that interacts with both the user and knowledge base to clarify how the user question relates to the stored information. MixAlign uses a language model to achieve automatic knowledge alignment and, if needed, further enhances this alignment through user clarifications. MixAlign focuses on utilizing grounding knowledge for faithful decision-making. In cases of uncertainty or unclear evidence, MixAlign generates a question seeking clarification from the user - a process referred to as human-assisted knowledge alignment.
Chain-of-Verification (CoVe): Dhuliawala et al. (2023) develop the CoVe method where the model
  1. 1.
    Drafts an initial response.
  2. 2.
    Plans verification questions to fact-check its draft.
  3. 3.
    Answers those questions independently so the answers are unbiased.
  4. 4.
    Generates a final verified response.
Experiments show CoVe decreases hallucinations across tasks like list-based Wikidata questions and long-form text generation. Given a user query, an LLM generates a baseline response that may contain inaccuracies like factual hallucinations. CoVe first generates verification questions to ask, then answers them to check for agreement.
Chain of Natural Language Inference (CoNLI):  Lei et al. (2023) address the challenge of hallucinations generated by LLMs when provided background context. Despite fluency in natural language generation, LLMs often produce ungrounded hallucinations unsupported by the given sources.
@@
The proposed hierarchical framework focuses on detecting and mitigating such hallucinations without requiring fine-tuning or domain-specific prompts. The framework utilizes Chain of Natural Language Inference (CoNLI) for state-of-the-art hallucination detection by identifying ungrounded content. Post-editing is then used to reduce hallucinations and enhance text quality without model adjustment. Extensive experiments on text-to-text datasets demonstrate effectiveness in both hallucination detection and reduction. By formulating detection as a chain of natural language inference tasks, the framework incorporates sentence and entity-level judgments with interpretability.
The plug-and-play framework allows seamless deployment across contexts with competitive hallucination detection and reduction performance while preserving text quality.
@@ 2.1.3 Prompt Tuning
SynTra: Large language models (LLMs) often exhibit hallucination in abstractive summarization tasks, even when the necessary information is present. Addressing this challenge is difficult due to the intricate evaluation of hallucination during optimization. Jones et al. (2023) introduce SynTra, a method that uses a synthetic task to efficiently reduce hallucination on downstream summarization tasks. SynTra optimizes the LLM’s system message via prefix-tuning on the synthetic task, then transfers this capability to more challenging, realistic summarization tasks. Experiments demonstrate reduced hallucination for two 13B parameter LLMs, highlighting the effectiveness of synthetic data for mitigating undesired behaviors.
@@ 3.3 Introducing faithfulness based loss function
Creating a metric to gauge how closely a model’s outputs match input data or ground truth is the task of this section. In this sense, faithfulness describes the model’s capacity to faithfully and properly reflect data from the input without adding errors, omissions, or distortions Chrysostomou and Aletras (2021). The following methods portray the use of technique:
Text Hallucination Mitigating (THAM) Framework: Yoon et al. (2022) introduce the THAM framework for Video-grounded Dialogue. THAM considers the text hallucination problem, which copies input texts for answer generation without the understanding of the question. It mitigates feature-level hallucination effects by introducing information-theoretic regularization. THAM framework incorporates Text Hallucination Regularization (THR) loss derived from the mutual information between the response language model and the proposed hallucination language model. Minimizing THR loss contributes to reducing indiscriminate text copying and boosting dialogue performances. THAM framework incorporates Text Hallucination Regularization loss derived from the proposed information-theoretic text hallucination measurement approach.
@@ 3.4 Supervised fine-tuning (SFT)
Knowledge Injection and Teacher-Student Approaches: Elaraby et al. (2023) focus on measuring and reducing hallucinations in weaker open-source large language models (LLMs) like BLOOM 7B Workshop et al. (2022). They introduce HALOCHECK, a lightweight knowledge-free framework to quantify hallucination severity in LLMs. The authors explore techniques like knowledge injection and teacher-student approaches to alleviate hallucinations in low-parameter LLMs. The framework uses sentence-level entailment to quantitatively assess hallucination levels.
The work aims to enhance smaller LLM knowledge through Knowledge Injection (KI) by fine-tuning with domain knowledge, without relying on expensive instructions from stronger models. They investigate leveraging a more powerful LLM like GPT-4 to guide weaker LLMs by generating detailed question answers. By assessing hallucination severity, they optimize teacher LLM engagement to reduce the computational costs of relying extensively on large models. This alleviates the need for frequent queries to the teacher model.
@@
Refusal-Aware Instruction Tuning (R-Tuning): In their recent work,  Zhang et al. (2023a) present a novel approach called R-Tuning for instilling refusal skills in large language models (LLMs). This approach formalizes the idea of identifying knowledge gaps between an LLM’s parametric knowledge and the instructional tuning data used to train it. Based on this knowledge gap, R-Tuning constructs refusal-aware training data to teach the LLM when to refrain from responding, specifically when a question falls outside its competence. The R-Tuning methodology involves two key steps:
@@
    Measuring the knowledge gap between the LLM’s parametric knowledge and the instructional tuning questions, to identify uncertain questions. By inferring on the training data once and comparing predictions to labels, the tuning data is separated into uncertain questions and certain questions.
@@
    Constructing refusal-aware training data by appending refusal expressions to uncertain training examples, before fine-tuning the LLM on this data.
Think While Effectively Articulating Knowledge (TWEAK): To reduce hallucinations, Qiu et al. (2023a) propose a new decoding method called TWEAK. The method treats the generated sequences at each step and their future sequences as hypotheses. It ranks each generation candidate based on how well their corresponding hypotheses support the input facts, using a Hypothesis Verification Model (HVM).
The authors tweak only the decoding process without retraining the generative models. This makes their approach easily integrated with any knowledge-to-text generator. Existing decoding methods like beam search sample candidates only based on predicted likelihood, without considering faithfulness. The authors propose a new dataset called FATE, which aligns input facts with original and counterfactual descriptions at the word level.